# Resilient Topology-Aware Coordination for Dynamic 3D UAV Networks under Node Failure

### v1

@@ 3.1.3. Inter-UAV Coordination and Control Plane
To enable cooperative behavior, the UAV swarm and the ground infrastructure are modeled as a dynamic graph $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t))$, where the vertex set $\mathcal{V}(t)$ incorporates both the surviving UAV nodes and the stationary GBS as defined in Section 3.1. An undirected edge exists between two operational UAVs if their Euclidean distance remains within the maximum communication range $R_{\mathrm{comm}}$. Distinctly, the GBS maintains persistent, long-range control-plane connectivity with all UAVs within the target area to facilitate global coordination and centralized value estimation.
@@
Through these coordination links, agents exchange low-overhead state information—including instantaneous velocities and latent policy embeddings—within their local one-hop neighborhood. This decentralized information exchange facilitates the proposed topology-aware attention mechanism, empowering the swarm to autonomously re-configure its coordination weights whenever $\mathcal{G}(t)$ undergoes topological ruptures due to sudden UAV node failures.
3.1.4. Wireless Backhaul (Data Plane)
@@
We assume a star topology for the data plane, where each surviving UAV $k\in\mathcal{V}(t)\setminus\{\text{GBS}\}$ maintains a high-capacity wireless backhaul link with the fixed GBS located at $\mathbf{p}_{\text{GBS}}$. The backhaul capacity is assumed to be sufficiently provisioned to support peak aggregate traffic. Consequently, the overall system performance is predominantly limited by the dynamic fronthaul access links between the UAVs and ground users rather than backhaul constraints.
3.2. User Distribution and Mobility Models
@@
The spatial dynamics of ground users are characterized through initialization and temporal evolution phases, both of which are contingent on the specific environmental scenario.
3.2.1. Initial Spatial Distribution
@@
At the beginning of an episode ($t=0$), user positions are initialized based on their logical topology to simulate heterogeneous service requirements.
@@
  • Clustered (Urban/Suburban): To represent high-density hotspots, users are partitioned into $K_{U}+1$ clusters using the K-Means algorithm, where $K_{U}$ denotes the number of available UAVs. The cluster whose centroid is closest to the GBS coordinates $\mathbf{p}_{\text{GBS}}$ is assigned to the terrestrial macro-cell for baseline service, while the remaining $K_{U}$ clusters are designated as target zones for the UAV swarm. Within each cluster, users follow a 2D Gaussian distribution $\mathcal{N}(\mathbf{c}_{j},\sigma^{2}_{c})$ around the respective centroid $\mathbf{c}_{j}$.
@@
  • Uniform (Rural): In sparse scenarios, the initial positions $\mathbf{q}_{m}(0)$ are uniformly distributed across the target area $\mathcal{D}$. In this case, users connect to either the GBS or a UAV based on the Max-RSSI association policy.
3.2.2. Mobility Dynamics
@@
We employ two distinct mobility models to capture realistic movement patterns. To avoid confusion with reinforcement learning parameters, we utilize specific notation for kinematic updates:
@@
  • Reference Point Group Mobility (RPGM) for Urban/Suburban: In high-density scenarios, users exhibit collective behavior. The motion of user $m$ belonging to cluster $j$ is governed by the reference velocity of the group center. The user position $\mathbf{q}_{m}(t)$ at time $t$ is expressed as $\mathbf{q}_{m}(t)=\mathbf{c}_{j}(t)+\mathbf{d}_{m,j}(t)$, where $\mathbf{c}_{j}(t)$ is the dynamic position of the group center and $\mathbf{d}_{m,j}(t)$ represents the relative displacement vector simulating individual deviation within the group.
@@
  • Gauss-Markov (GM) for Rural: For independent users, we minimize abrupt velocity changes through the GM model. Let $\mathbf{v}_{m}(t)$ denote the velocity of user $m$. The update follows the rule:
    (1) $\mathbf{v}_{m}(t+1)=\alpha_{\text{GM}}\mathbf{v}_{m}(t)+(1-\alpha_{\text{GM}})\bar{\mathbf{v}}_{m}+\sqrt{1-\alpha_{\text{GM}}^{2}}\boldsymbol{\epsilon}_{m}(t),$
@@
    where $\alpha_{\text{GM}}\in[0,1]$ is the memory parameter controlling temporal correlation. The term $\bar{\mathbf{v}}_{m}$ is the asymptotic mean velocity and $\boldsymbol{\epsilon}_{m}(t)$ is an independent Gaussian noise process.
3.3. Channel Propagation and Interference Models
@@
The communication performance is governed by the propagation characteristics of both terrestrial and air-to-ground links.
3.3.1. Terrestrial Channel (GBS-User)
@@
The path loss $L_{\text{GBS},m}(t)$ in dB between the macro GBS and user $m$ is modeled using a log-normal shadowing approach:
(2) $L_{\text{GBS},m}(t)=\text{PL}(d_{0})+10\kappa\log_{10}\left(\frac{d_{\text{GBS},m}(t)}{d_{0}}\right)+\chi_{\sigma},$
@@
where $\text{PL}(d_{0})$ represents the path loss at a reference distance $d_{0}$, $\kappa$ is the path loss exponent, and $d_{\text{GBS},m}(t)$ denotes the 3D Euclidean distance. The term $\chi_{\sigma}$ accounts for shadow fading, which follows a zero-mean Gaussian distribution with a standard deviation $\sigma$.
3.3.2. Air-to-Ground Channel (UAV-User)
@@
The link between an operational UAV $k\in\mathcal{V}(t)$ and user $m$ incorporates probabilistic Line-of-Sight (LoS) propagation. The LoS probability $P_{\mathrm{LoS},k,m}(t)$ is calculated as a function of the elevation angle between the UAV and the user. Consequently, the average path loss $\overline{\text{PL}}_{k,m}(t)$ is expressed as follows:
@@
(3) $\overline{\text{PL}}_{k,m}(t)=P_{\mathrm{LoS},k,m}(t)\eta_{\mathrm{LoS}}+(1-P_{\mathrm{LoS},k,m}(t))\eta_{\mathrm{NLoS}}+20\log_{10}\left(\frac{4\pi f_{c}d_{k,m}(t)}{c}\right),$
@@
where $\eta_{\mathrm{LoS}}$ and $\eta_{\mathrm{NLoS}}$ are the excessive path loss coefficients for LoS and NLoS conditions, $f_{c}$ is the carrier frequency, and $c$ is the speed of light.
3.3.3. Directional Antenna Gain
@@
Each UAV is equipped with a directional antenna to suppress co-channel interference. Let $\phi_{k,m}(t)$ be the off-axis angle of user $m$ relative to the vertical boresight of UAV $k$. The antenna gain $G_{\mathrm{tx}}(\phi_{k,m})$ follows a flat-top model:
@@
(4) $G_{\mathrm{tx}}(\phi_{k,m})=\begin{cases}G_{\mathrm{main}},&\text{if }|\phi_{k,m}(t)|\leq\Theta_{B}\\ G_{\mathrm{side}},&\text{otherwise}\end{cases},$
@@
where $G_{\mathrm{main}}$ and $G_{\mathrm{side}}$ are the gains for the main-lobe and side-lobe, respectively, and $2\Theta_{B}$ represents the half-power beamwidth. The effective channel gain is defined as $h_{k,m}(t)=G_{\mathrm{tx}}(\phi_{k,m})\cdot 10^{-\overline{\text{PL}}_{k,m}(t)/10}$.
3.4. Interference Model and Data Rate
@@
We assume a universal frequency reuse strategy where all operational UAVs in the set $\mathcal{V}(t)$ share the same frequency band $B$.
3.4.1. Signal-to-Interference-plus-Noise Ratio (SINR)
@@
Assuming a Max-RSSI association policy where user $m$ connects to an active UAV $k\in\mathcal{V}(t)\setminus\{\text{GBS}\}$, the received SINR $\gamma_{k,m}(t)$ is formulated as follows:
(5) $\gamma_{k,m}(t)=\frac{P_{\text{tx}}h_{k,m}(t)}{N_{0}B+\sum_{j\in\mathcal{V}(t)\setminus{k,\text{GBS}}}P_{\text{tx}}h_{j,m}(t)+P_{\text{GBS}}h_{\text{GBS},m}(t)},$
@@
where $P_{\text{GBS}}h_{\text{GBS},m}(t)$ represents the co-channel interference from the terrestrial GBS. The summation in the denominator is restricted to the active UAV set $\mathcal{V}(t)\setminus\{k,\text{GBS}\}$, which indicates that failed nodes do not contribute to co-channel interference. This formulation, combined with the persistent interference from the GBS, ensures that aggregate interference is calculated only over the surviving serving nodes.
3.4.2. Achievable Data Rate
@@
Based on the assumption of equal bandwidth sharing among users associated with the same UAV, the downlink data rate $R_{m}(t)$ for user $m$ served by UAV $k$ is given by
(6) $R_{m}(t)=\frac{B}{N_{k}(t)}\log_{2}(1+\gamma_{k,m}(t)),$
@@
where $N_{k}(t)$ denotes the total load representing the number of users currently served by UAV $k$.
3.5. UAV Power Consumption Model
@@
We adopt an analytical power model for rotary-wing UAVs to characterize the energy efficiency of the networks. Let $v_{k}(t)=\|\mathbf{v}_{k}(t)\|$ be the instantaneous scalar speed of UAV $k$. The propulsion power $P_{k}^{\text{fly}}(t)$ required for maintaining flight at speed $v_{k}(t)$ is formulated as follows (Zeng et al., 2019):
@@
(7) $P_{k}^{\text{fly}}(t)=P_{0}\left(1+\frac{3v_{k}(t)^{2}}{U_{\text{tip}}^{2}}\right)+P_{i}\left(\sqrt{1+\frac{v_{k}(t)^{4}}{4v_{0}^{4}}}-\frac{v_{k}(t)^{2}}{2v_{0}^{2}}\right)^{1/2}+\frac{1}{2}d_{\text{fuse}}\rho sAv_{k}(t)^{3},$
@@
where $P_{0}$ and $P_{i}$ represent the blade profile power and induced power in hovering status, respectively. The constant $U_{\text{tip}}$ denotes the tip speed of the rotor blade, and $v_{0}$ is the mean rotor induced velocity when hovering. The parameters $d_{\text{fuse}}$, $\rho$, $s$, and $A$ represent the fuselage drag ratio, air density, rotor solidity, and rotor disc area, respectively. Specifically, the three terms in (7) correspond to the blade profile power, induced power, and parasitic power required to overcome fuselage drag.
@@
Beyond the propulsion of aerial nodes, the total network power consumption $P_{\text{total}}(t)$ incorporates the energy cost of both the dynamic UAV swarm and the fixed ground infrastructure. It is expressed as:
(8) $P_{\text{total}}(t)=\sum_{k=1}^{K_{U}}\xi_{k}(t)(P_{k}^{\text{fly}}(t)+P_{\text{comm}})+P_{\text{GBS}},$
@@
where $K_{U}=4$ is the number of mobile UAVs, and the total serving nodes are $K=K_{U}+1$. $P_{\text{GBS}}$ denotes the constant power consumption of the stationary GBS, which includes its signal processing and backhaul transmission circuitry. The inclusion of the survival indicator $\xi_{k}(t)$ ensures that power is only aggregated over operational UAVs, reflecting the physical reality that a failed node no longer draws propulsion or communication power from the system’s aggregate resources.
4. Problem Formulation
4.1. Network Utility Maximization
@@
The primary objective is to maintain service continuity, load balance, and system stability amidst rapid user mobility and potential node failures. We define a composite utility function $U(t)$ derived from key performance indicators (KPIs) reflecting both Quality of Service (QoS) and Quality of Experience (QoE).
4.1.1. System Capacity
@@
4.1.2. Global Energy Efficiency

### v2

@@ III-A3 Inter-UAV Coordination and Control Plane
To enable cooperative behavior, the UAV swarm and ground infrastructure are modeled as a dynamic graph $\mathcal{G}(t)=(\mathcal{V}(t),\mathcal{E}(t))$, where $\mathcal{V}(t)$ incorporates surviving UAV nodes and the stationary GBS. An undirected edge exists between operational UAVs if their distance is within the maximum communication range $R_{\mathrm{comm}}$. The GBS maintains persistent, long-range control-plane connectivity with all UAVs to facilitate global coordination and centralized value estimation.
@@
Through these coordination links, agents exchange low-overhead state information, including instantaneous velocities and latent policy embeddings, within their local one-hop neighborhood. This decentralized exchange facilitates the proposed topology-aware attention mechanism, empowering the swarm to autonomously re-configure coordination weights whenever $\mathcal{G}(t)$ undergoes topological ruptures due to node failures.
Figure 1: Illustration of the dynamic 3D aerial-ground integrated network scenario. A fleet of UAVs functions as flying base stations to assist the Macro GBS in serving high-mobility vehicular platoons at a traffic intersection. The red solid arrows represent the high-capacity wireless backhaul links connected to the GBS (Data Plane), forming a star topology. The blue dotted lines denote the inter-UAV coordination links for exchanging local state information (Control Plane). The dashed ovals indicate the directional service footprints dynamically tracking the moving user clusters.
III-A4 Wireless Backhaul (Data Plane)
@@
We assume a star topology for the data plane, where each surviving UAV $k\in\mathcal{V}(t)\setminus\{\text{GBS}\}$ maintains a high-capacity wireless backhaul link with the GBS at $\mathbf{p}_{\text{GBS}}$. The backhaul capacity is assumed to be sufficiently provisioned to support peak aggregate traffic. Consequently, system performance is limited by the dynamic fronthaul access links between UAVs and ground users rather than backhaul constraints.
III-B User Distribution and Mobility Models
@@
Ground user spatial dynamics involve initialization and temporal evolution phases contingent on the environmental scenario.
III-B1 Initial Spatial Distribution
@@
At $t=0$, user positions are initialized based on logical topology to simulate heterogeneous service requirements.
@@
  • Clustered (Urban/Suburban): To represent hotspots, users are partitioned into $K_{U}+1$ clusters via K-Means. The cluster with the centroid closest to $\mathbf{p}_{\text{GBS}}$ is assigned to the terrestrial macro-cell, while the remaining $K_{U}$ clusters are designated as UAV target zones. Within each cluster, users follow a 2D Gaussian distribution $\mathcal{N}(\mathbf{c}_{j},\sigma^{2}_{c})$ around centroid $\mathbf{c}_{j}$.
@@
  • Uniform (Rural): In sparse scenarios, initial positions $\mathbf{q}_{m}(0)$ are uniformly distributed across the target area $\mathcal{D}$. Users connect to either the GBS or a UAV based on the Max-RSSI association policy.
III-B2 Mobility Dynamics
@@
We employ two mobility models to capture realistic movement patterns using specific kinematic notation.
@@
  • Reference Point Group Mobility (RPGM): In high-density scenarios, user motion follows collective behavior. The position $\mathbf{q}_{m}(t)$ of user $m$ in cluster $j$ is $\mathbf{q}_{m}(t)=\mathbf{c}_{j}(t)+\mathbf{d}_{m,j}(t)$, where $\mathbf{c}_{j}(t)$ is the dynamic group center and $\mathbf{d}_{m,j}(t)$ represents a relative displacement vector simulating individual deviation.
@@
  • Gauss-Markov (GM): For independent users, we minimize abrupt velocity changes through the GM model. Let $\mathbf{v}_{m}(t)$ denote the velocity of user $m$. The update follows:
    $\mathbf{v}_{m}(t+1)=\alpha_{\text{GM}}\mathbf{v}_{m}(t)+(1-\alpha_{\text{GM}})\bar{\mathbf{v}}_{m}+\sqrt{1-\alpha_{\text{GM}}^{2}}\bm{\epsilon}_{m}(t),$ (1)
@@
    where $\alpha_{\text{GM}}\in[0,1]$ is the memory parameter and $\bar{\mathbf{v}}_{m}$ is the asymptotic mean velocity. The term $\bm{\epsilon}_{m}(t)$ represents an independent Gaussian noise process.
III-C Channel Propagation Model
@@
Communication performance is governed by the propagation characteristics of terrestrial and air-to-ground links.
III-C1 Terrestrial Channel (GBS-User)
@@
The path loss $L_{\text{GBS},m}(t)$ in dB between the GBS and user $m$ is modeled via log-normal shadowing:
$L_{\text{GBS},m}(t)=\text{PL}(d_{0})+10\kappa\log_{10}\left(\frac{d_{\text{GBS},m}(t)}{d_{0}}\right)+\chi_{\sigma},$ (2)
@@
where $\text{PL}(d_{0})$ is the path loss at reference distance $d_{0}$, $\kappa$ is the path loss exponent, and $d_{\text{GBS},m}(t)$ is the 3D Euclidean distance. The term $\chi_{\sigma}$ accounts for shadow fading following a zero-mean Gaussian distribution with standard deviation $\sigma$.
III-C2 Air-to-Ground Channel (UAV-User)
@@
The link between an operational UAV $k\in\mathcal{V}(t)$ and user $m$ incorporates probabilistic Line-of-Sight (LoS) propagation. The LoS probability $P_{\mathrm{LoS},k,m}(t)$ depends on the elevation angle. The average path loss $\overline{\text{PL}}_{k,m}(t)$ is:
@@
$\begin{split}\overline{\text{PL}}_{k,m}(t)=&P_{\mathrm{LoS},k,m}(t)\eta_{\mathrm{LoS}}+(1-P_{\mathrm{LoS},k,m}(t))\eta_{\mathrm{NLoS}}\\ &+20\log_{10}\left(\frac{4\pi f_{c}d_{k,m}(t)}{c}\right),\end{split}$ (3)
@@
where $\eta_{\mathrm{LoS}}$ and $\eta_{\mathrm{NLoS}}$ are excessive path loss coefficients for LoS and NLoS conditions, $f_{c}$ is the carrier frequency, and $c$ is the speed of light.
III-C3 Directional Antenna Gain
@@
Each UAV uses a directional antenna to suppress co-channel interference. Let $\phi_{k,m}(t)$ be the off-axis angle of user $m$ relative to the boresight of UAV $k$. The antenna gain $G_{\mathrm{tx}}(\phi_{k,m})$ follows a flat-top model:
@@
$G_{\mathrm{tx}}(\phi_{k,m})=\begin{cases}G_{\mathrm{main}},&\text{if }|\phi_{k,m}(t)|\leq\Theta_{B}\\ G_{\mathrm{side}},&\text{otherwise}\end{cases},$ (4)
@@
where $G_{\mathrm{main}}$ and $G_{\mathrm{side}}$ are the main-lobe and side-lobe gains, and $2\Theta_{B}$ is the beamwidth. The effective channel gain is $h_{k,m}(t)=G_{\mathrm{tx}}(\phi_{k,m})\cdot 10^{-\overline{\text{PL}}_{k,m}(t)/10}$.
III-D Interference Model and Data Rate
@@
We assume a universal frequency reuse strategy where all operational UAVs in $\mathcal{V}(t)$ share the same frequency band $B$.
III-D1 Signal-to-Interference-plus-Noise Ratio (SINR)
@@
Under the Max-RSSI association policy where user $m$ connects to an active UAV $k\in\mathcal{V}(t)\setminus\{\text{GBS}\}$, the received SINR $\gamma_{k,m}(t)$ is:
$\gamma_{k,m}(t)=\frac{P_{\text{tx}}h_{k,m}(t)}{N_{0}B+\sum_{j\in\mathcal{V}(t)\setminus{k,\text{GBS}}}P_{\text{tx}}h_{j,m}(t)+P_{\text{GBS}}h_{\text{GBS},m}(t)},$ (5)
@@
where $P_{\text{GBS}}h_{\text{GBS},m}(t)$ represents co-channel interference from the terrestrial GBS. The summation in the denominator is restricted to the active set $\mathcal{V}(t)\setminus\{k,\text{GBS}\}$, ensuring that failed nodes do not contribute to the interference environment. Combined with persistent GBS interference, this formulation calculates aggregate interference only over surviving nodes.
III-D2 Achievable Data Rate
@@
Assuming equal bandwidth sharing among users associated with the same UAV, the downlink data rate $R_{m}(t)$ for user $m$ served by UAV $k$ is:
$R_{m}(t)=\frac{B}{N_{k}(t)}\log_{2}(1+\gamma_{k,m}(t)),$ (6)
@@
where $N_{k}(t)$ denotes the total load, representing the number of users currently served by UAV $k$.
III-E UAV Power Consumption Model
@@
We adopt an analytical power model for rotary-wing UAVs to evaluate network energy efficiency. Let $v_{k}(t)=\|\mathbf{v}_{k}(t)\|$ be the scalar speed of UAV $k$. The propulsion power $P_{k}^{\text{fly}}(t)$ required to maintain flight at speed $v_{k}(t)$ is formulated as [30]:
@@
$\begin{split}P_{k}^{\text{fly}}(t)=&P_{0}\left(1+\frac{3v_{k}(t)^{2}}{U_{\text{tip}}^{2}}\right)+\frac{1}{2}d_{\text{fuse}}\rho sAv_{k}(t)^{3}\\ &+P_{i}\left(\sqrt{1+\frac{v_{k}(t)^{4}}{4v_{0}^{4}}}-\frac{v_{k}(t)^{2}}{2v_{0}^{2}}\right)^{1/2},\end{split}$ (7)
@@
where $P_{0}$ and $P_{i}$ are the blade profile and induced power in hovering, respectively. $U_{\text{tip}}$ is the rotor blade tip speed, and $v_{0}$ is the mean rotor induced velocity when hovering. The parameters $d_{\text{fuse}}$, $\rho$, $s$, and $A$ denote fuselage drag ratio, air density, rotor solidity, and rotor disc area. The three terms in (7) correspond to blade profile, induced, and parasitic power.
@@
Total network power consumption $P_{\text{total}}(t)$ incorporates the energy cost of the UAV swarm and the ground infrastructure:
$P_{\text{total}}(t)=\sum_{k=1}^{K_{U}}\xi_{k}(t)(P_{k}^{\text{fly}}(t)+P_{\text{comm}})+P_{\text{GBS}},$ (8)
@@
where $K_{U}=4$ and $K=K_{U}+1$. $P_{\text{GBS}}$ denotes the constant power of the stationary GBS, including signal processing and backhaul circuitry. The survival indicator $\xi_{k}(t)$ ensures power is aggregated only over operational UAVs, as failed nodes do not draw propulsion or communication power.
IV Problem Formulation
IV-A Network Utility Maximization
@@
The primary objective is to maintain service continuity and system stability amidst rapid user mobility and potential node failures. We define a composite utility function $U(t)$ derived from key performance indicators (KPIs) reflecting both Quality of Service (QoS) and Quality of Experience (QoE).
IV-A1 System Capacity
@@
IV-A2 Global Energy Efficiency